OpenStudy (anonymous):

@astrophysics If the side lengths of a cube are 14 feet, what is the correct way to write the expression to represent the volume of the cube in exponential form? 143 314 14 ⋅ 3 14 ⋅ 14 ⋅ 14

OpenStudy (astrophysics):

This seems fun, ok so you know that \[y^2 = y \times y\] \[y^5 = y \times y \times y \times y \times y\] right? The exponent just means you multiply it by itself that many times, does that make sense so far?

OpenStudy (anonymous):

a little bit

OpenStudy (astrophysics):

Ok, so when we talk about areas we have 2 dimensions that would mean we have something squared, when we talk about volume we have 3 dimensions so we cube it, let me draw what I mean.  so A here represents area and we know to get area of a square we write it as \[a^2 = a \times a~~~~14^2 = 14 \times 14\] now we have a cube and to find it's volume, we deal with 3 dimensions so it looks as such  notice all the dimensions in a cube are the same so we have volume as \[V =a^3\] so what can we conclude our answer to be? (I know this might be a lot at once, but I hope it helps you understand)